FOURIER TRANSFORM MICROWAVE SPECTRUM OF AlCCH (X^{1}Σ^{+}) AND ITS ^{13}C/D ISOTOPO-LOGUES

M. SUN, Department of Chemistry, Department of Astronomy, and Steward Observatory, University of Arizona, Tucson, AZ 85721; D. J. CLOUTHIER, Department of Chemistry, University of Kentucky, Lexington, KY 40506; and \underline{L. M. ZIURYS}, Department of Chemistry, Department of Astronomy, and Steward Observatory, University of Arizona, Tucson, AZ 85721.

The microwave spectrum of AlCCH (X^{1}Σ^{+}) has been measured using Fourier transform microwave (FTMW) techniques in the frequency range of 9-40 GHz. This molecule was created in a supersonic expansion by the reaction of trimethyl aluminum, Al(CH_{3})_{3}, and acetylene, C_{2}H_{2}, or methane, CH_{4}, diluted in argon carrier gas, using a pulsed nozzle coupled with a dc discharge. Spectra of Al^{12}C^{13}CH, Al^{13}C^{12}CH, Al^{13}C^{13}CH, Al^{12}C^{12}CD have also been recorded. From these data, rotational constants and Al and D hyperfine parameters have been determined, as well as the r_{m}^{(1)} structure. AlCCH is a model system for heteroatom dicarbide species.